# Exploration of Quantum Computing in Materials Discovery for Direct Air Capture Applications

### v1

@@ Abstract
Direct air capture is a promising method for mitigating climate change. Solid sorbents, such as metal-organic frameworks, have been considered for this task, but their potential for deployment at scale has not been fully realized. The computational discovery of sorbent materials is a daunting task, given the vast search space, and the fact that their real-world performance depends on their ability to selectively bind CO2 molecules while avoiding other more abundant flue gas components. Quantum computing can potentially accelerate the discovery of novel materials for direct air capture as an alternative way to compute binding energies. In this work, we demonstrate methods and algorithms that enable these calculations in current quantum computing devices. We simulate the potential energy surfaces of CO2, N2, and H2O molecules at the Mg+2metal center that represents the binding sites of typical metal-organic frameworks. We apply the qubit-ADAPT-VQE technique to run simulations on both classical and quantum hardware, and achieve reasonable accuracy while maintaining hardware efficiency, even when compared to more established methods such as UCCSD-VQE.
@@ 1 Introduction
Technologies for capturing greenhouse gases (GHG) are essential climate mitigation tools that can help us reduce GHG atmospheric concentration and limit global warming.1 Existing energy generation infrastructure, such as natural gas and coal-fired power plants, can be retrofitted to perform post-combustion capture of carbon dioxide. This strategy can extend the lifetime of existing infrastructure and minimize the impact of energy generation to our global carbon budget. Nevertheless, merely bringing emissions to zero is not sufficient for achieving our climate goals. We need to actively remove carbon dioxide from the atmosphere if we want to limit the temperature increase to manageable levels. Direct air capture (DAC) is an emerging technology that employs sorbents to capture CO2 from the atmosphere.
@@
Liquid sorbent technologies, such as amine scrubbing, are the most well-established for capturing carbon dioxide from the air. Nevertheless, despite its fast adsorption rate and high CO2 saturation, amines present disadvantages that hinder their widespread implementation: they are corrosive, and due to their high desorption energy requirement, are responsible for 60-80% of total operating costs of CCS.2 Solid sorbent materials for carbon capture offer an alternative approach to overcome the limitations of amine-based (liquid) adsorption. This material class includes a diverse range of nanoporous solids, such as zeolites, metal-organic frameworks, covalent-organic frameworks, zeolitic imidazolate frameworks and porous polymer networks, among others. Porous solid sorbents can selectively adsorb CO2 without bond formation (i.e., via physisorption), which drastically lowers the regeneration energy and, therefore, the operating costs. Another advantage over amines is that the amount of CO2 adsorbed is not proportional to the amount of solid material, given that the adsorption takes place at the void space (e.g., cages, channels, or slits) that exists inside the solid. On the other hand, solid sorbents can be very sensitive to the presence of water vapor in the flue gas, leading to decreased selectivity due to competitive adsorption with CO2 and incomplete regeneration due to water passivation of adsorption sites.
@@
Among the many solid materials classes considered, Metal-Organic Frameworks (MOFs) have attracted considerable attention due to their customizable porous structure and chemical diversity.3 They are composed of two types of building blocks – i.e., organic linkers and metal ions – which can be combined to form nanoporous structures with channels whose diameters range from a few to tens of Angstroms. The chemical and geometrical aspects can be leveraged independently to optimize both the absolute (adsorption capacity) and relative (adsorption selectivity) metrics that determine what is a good carbon capture material. The ample variety of MOFs, one of its main advantages, is also the origin of significant computational challenges for the selection of the best candidate material for a given application. There are databases of hypothetical MOFs,4, 5 as well as experimentally verified MOFs,6 amounting to $10^{5}-10^{6}$ materials in total.
@@
In this work, we perform Quantum Chemistry calculations, on classical and on quantum hardware, to estimate the binding potential of a few molecules of interest present in the atmosphere (CO2, H2O and N2) to simplified model systems for carbon capture. We simplified the search space for a carbon capture material by focusing on the X-MOF-74 family (where X=Mg+2, Ni+2, Fe+2, Co+2, Zn+2, Mn+2, Cu+2) as a template for exploring the chemical diversity in MOFs. The MOF-74 family is known for the presence of coordinatively unsaturated metal ions (open metal sites) that can form a bond with adsorbate molecules, greatly influencing the adsorption strength and selectivity, especially at low pressures. The formation of chemical bonds with the adsorbate molecule is a phenomenon known as chemisorption and determines, to a great extent, the desorption energy cost.
@@ 1.1.1 Species selection
Given the relevance of these metal ions to the application in MOFs, we chose to focus on Mg+2metal ion and its pairings with all the three provided gas molecules (H2O, CO2, and N2), amounting to six distinct systems. In short, Mg+2 + H2O should be the simplest system as it contains the least number of electrons.
@@ 1.1.2 Simulation methods
We chose PySCF,7, 8 as the driver for performing classical benchmarks. We ran three types of classical benchmarks with PySCF: Hartree-Fock (HF) and Density Functional Theory (DFT). The chosen basis set was 6-31g* and was kept consistent across all calculations, including for the quantum calculations. For DFT we chose the B3LYP functional. This particular choice of parameters was motivated by the methodology employed to compute 3D structures (SD-file) of the H2O, CO2, and N2 molecules found in the NIST Chemistry WebBook, from which we also took the molecular structures used in the calculations. 9
1.1.3 Considerations for Quantum devices
@@
  • A usual VQE calculation with an application-specific, compute-heavy UCCSD ansatz,14 to be treated as the quantum benchmark, as it provides the best possible result for this kind of application. Unfortunately, it is challenging to use this approach on current quantum devices due to its complexity.
  • A usual VQE calculation with a general-purpose, hardware-efficient heuristic ansatz that can run on current quantum devices. In our case, we determined that an ansatz with RY and RZ rotations, which we call RYRZ, converges to the intended result despite requiring more iterations.
  • An implementation of the ADAPT-VQE algorithm15 that utilizes qubit operators in the operator pool. This method is called qubit-ADAPT-VQE and has the advantage of giving similar results to UCCSD while having much lower requirements on the quantum hardware.16
@@ 1.1.4 Active Space Selection
To limit the computational cost for the VQE calculations, we performed a suitable active space selection. We used a known method implemented within PySCF called atomic valence active space (AVAS).17 AVAS is an automated method to construct an active space for electronic structure calculations in strongly correlated chemical systems, where single-determinant wave functions are inadequate. It focuses on identifying specific valence atomic orbitals, such as transition metal d orbitals or those involved in bond dissociation, as a basis for defining the active space. The approach leverages density matrix embedding theory (DMET) to determine the minimal set of orbitals required to capture the essential electronic features, aiming for an active space no more than twice the size of the initially chosen orbitals. This method does not differentiate between “fragment” and “bath” orbitals, instead maintaining the occupied and virtual character of active orbitals, facilitating a natural truncation process to minimize active space size while capturing critical determinants.
@@ 2.1 Convergence studies
The qubit-ADAPT-VQE ansatz accomplishes fast convergence while remaining hardware efficient, as we can see from Table 2 comparing the sizes of the circuits between methods with the largest circuit that qubit-ADAPT-VQE uses. These circuit sizes are expressed in terms of the gates that IBM quantum devices execute and were built using the transpilation procedure from Qiskit. The most important metric here is the number of CNOT gates as those have error rates that are orders of magnitude higher than the single-qubit gates.
Table 2: Comparing the circuit sizes generated by different ansatze. qubit-ADAPT-VQE manages to get satisfactory results while using the shallowest circuits. Circuits are described in terms of the number of gate operations and depth required to run on IBM quantum devices.
@@ 2.2 Mg+2 systems
Due to the excellent cost-benefit obtained from the qubit-ADAPT-VQE ansatz in the simulations, we selected this as our best candidate to run experiments on the quantum hardware. However, to run these calculations on the available quantum hardware, we had to further reduce our active space to three electrons and four molecular orbitals (3,4), which only required 4 qubits. Despite nominally supporting higher qubit counts, current devices are not suitable for reliably running circuits of the required depth for this application due to a lack of fault tolerance. Again, the molecular orbitals were chosen around the HOMO-LUMO level and ordered by energy. We set the limit on the number of shots to 1024, which takes from five to eight hours depending on the separation distance.
The 4-qubit active space limitation led us to focus only on the Mg+2 systems for hardware experiments. This active space corresponds to $\approx 5\%$ of the energy contributions to the ground state, therefore we saw no point in simulating systems with a higher electron count as it would mean an even lower representation of the results obtained. We used Qiskit Runtime and ibmq_montreal to run all our quantum hardware experiments, and we employed error mitigation techniques such as Twirled Readout Error Extinction (T-REx).19.
The PES was built by computing the energy for different separations between metal ions and molecules along the bond axis. For all simulations and experiments, we evaluate the energy at distances between 0.3 and 3.6 Å, in increments of 0.3 Å.
@@
We simulated Mg+2 binding with all of the suggested gas molecules using the three VQE methods mentioned before: RYRZ-VQE, UCCSD-VQE, and qubit-ADAPT-VQE, on a (6,7) active space. We evaluate their accuracy by comparing their PES to that obtained via direct diagonalization of the Hamiltonian with NumPy.
@@
Based on the accuracy of qubit-ADAPT-VQE results in Figure 4, and the superior performance over UCCSD-VQE demonstrated in Figure 3, Table 1 and Table 2, we confirm the choice of qubit-ADAPT-VQE for running VQE on quantum hardware. We limit ourselves to the smallest possible system, namely, Mg+2 + H2O for the reasons explained previously. We ran qubit-ADAPT-VQE on quantum hardware with an active space of (3,4) to comply with the limitations of current quantum devices.
The results, shown in Figure 5, also provide a comparison to a quantum hardware calculation. The Morse potential fit to the hardware result matches quite closely those of the direct diagonalization via NumPy and simulated VQE, despite the level of noise that is inherent to calculations on current quantum devices. In this hardware experiment, we used 1024 shots and employed the T-REx noise mitigation technique which proved sufficient to handle the noise profile of the quantum device. Increasing the number of shots did not improve the accuracy of the experimental values, and using more robust error mitigation strategies, such as Zero Noise Extrapolation, would increase the already large variance in the results.
@@
We extracted observables such as the equilibrium distance and the binding energy from the adjusted Morse potentials to the Mg+2 + H2O PES curves. Although we cannot claim to have achieved agreement to within chemical accuracy ($\sim 1\ \text{mHa}$), on Table 3 we observe an excellent match between qubit-ADAPT-VQE and NumPy, and a reasonably good match with ibmq_montreal, considering the higher uncertainties caused by the quantum device noise.
@@
We compare the potential energy surfaces computed using classical methods from PySCF, such as HF and DFT, to the best-in-class simulated VQE method, namely, qubit-ADAPT-VQE in a (6,7) active space. In Figure 6, we show the agreement between quantum methods and their classical benchmarks. In particular, for Mg+2 + H2O, we also include in Figure 6(a) the comparison to the quantum hardware experiments running qubit-ADAPT-VQE with ibmq_montreal in a (3,4) active space.
Overall, the simulated qubit-ADAPT-VQE results match very well the HF benchmark for all Mg+2-based systems studied, with the DFT benchmark coming quite close. The PES obtained by running in the quantum hardware, despite its inherent noise, is also in good agreement with the classical benchmarks. This conclusion is supported not only by the visual inspection of Figure 6 but also by a comparison of the numeric results of the observables extracted from the PES fit in Table 4. Here, the Binding Energy values all agree within the numeric uncertainty, and while the equilibrium distances between qubit-ADAPT-VQE and HF concur, they are found to be 2-3% higher than what is predicted by DFT.
It’s worth noting the high uncertainties in the hardware results. This comes as a cost of using error mitigation techniques. while accuracy is increased there is a loss in precision. Considering its already relatively high value we opted not to use more robust error mitigation techniques that would increase it further.
@@
Table 4: Comparison of Equilibrium Distances and Binding Energies for Mg+2-based systems obtained from the fitted Morse potential across different methods, both quantum (simulations and experiments) and classical. The Binding Energy values all agree to within the numeric uncertainty. The equilibrium distances agree between qubit-ADAPT-VQE and HF, but are 2-3% higher than what is predicted by DFT.
@@ 2.3 Conclusions
The authors thank Max Rossmannek (IBM Research – Zurich) for providing help in implementing qubit-ADAPT-VQE and the broader Qiskit community for giving support via the Slack channel and GitHub pages. We thank Vidushi Sharma (IBM Research – Almaden) for her support in the interpretation of the classical benchmark methods. We thank Felipe Lopes de Oliveira (IBM Research – Brazil) for suggesting the Morse potential fit as a means to extract the observables. We thank Ivan S. Oliveira (CBPF) for insightful discussions on quantum computing. The authors would like to acknowledge the support of Alexandre Pfeifer and Bruno Flach (IBM Research - Brazil) for the realisation of this work.

### v2

@@ Abstract
Direct air capture (DAC) of carbon dioxide is a promising method for mitigating climate change. Solid sorbents, such as metal-organic frameworks, are currently being tested for DAC application. However, their potential for deployment at scale has not been fully realized. The computational discovery of solid sorbents is challenging, given the vast chemical search space and the DAC requirements for molecular selectivity. Quantum computing can potentially accelerate the discovery of solid sorbents for DAC by predicting molecular binding energies. In this work, we explore simulation methods and algorithms for predicting gas adsorption in metal-organic frameworks using a quantum computer. Specifically, we simulate the potential energy surfaces of CO2, N2, and H2O molecules at the Mg+2 metal center that represents the binding sites of typical metal-organic frameworks. We apply the qubit-ADAPT-VQE technique to run simulations on both classical computing and quantum computing hardware, and achieve reasonable accuracy while maintaining hardware efficiency.
@@ 1 Introduction
Technologies for capturing greenhouse gases (GHG) are essential climate mitigation tools that can help reduce GHG atmospheric concentration and limit global warming.1 Existing energy generation infrastructure, such as natural gas and coal-fired power plants, can be retrofitted to perform post-combustion capture of carbon dioxide. This strategy can extend the lifetime of existing infrastructure and minimize the impact of energy generation to the global carbon budget. Nevertheless, merely bringing emissions to zero is not sufficient for achieving climate goals. Carbon dioxide removal from the atmosphere is needed for limiting the temperature increase international2022direct. Direct air capture (DAC) is an emerging carbon removal technology that employs sorbents to capture CO2 directly from the atmosphere.
@@
Liquid sorbent based technologies, such as amine scrubbing, are well-established for capturing carbon dioxide.sanz2016direct Nevertheless, despite their fast adsorption rate and high CO2 saturation, amines have certain disadvantages. They are corrosive, and due to their high desorption energy requirement, are responsible for 60-80% of the total operating costs of CCS.2, sabatino2021comparative Solid sorbent materials for carbon dioxide capture offer an alternative approach to overcome the limitations of amine-based, liquid adsorption. This material class includes a diverse range of nanoporous solids, such as zeolites, metal-organic frameworks, covalent-organic frameworks, zeolitic imidazolate frameworks and porous polymer networks, among others. Porous solid sorbents can selectively adsorb CO2 without bond formation (i.e., via physisorption), which drastically lowers the regeneration energy and, therefore, the operating costs.kumar2015direct On the other hand, solid sorbents can be very sensitive to the presence of water vapor in the atmosphere, leading to decreased selectivity due to competitive adsorption with CO2 and incomplete regeneration due to water passivation of adsorption sites.decoste2013effect
@@
Among the solid materials currently under investigation, Metal-Organic Frameworks (MOFs) have attracted attention due to their customizable porous structure and large chemical diversity.3 MOFs are composed of two types of building blocks, i.e., organic linkers and metal ions, which can be combined to form nanoporous structures with channels whose diameters range from a few to tens of Angstroms. Their chemical and geometrical aspects can be modified, to a large extent, independently for optimizing both the absolute (adsorption capacity) and relative (adsorption selectivity) metrics that determine capture and separation performance.yao2021inverse The large variety of MOFs poses significant computational challenges for the selection of the best candidate material for a given application. There are databases of hypothetical MOFs,4, 5 as well as experimentally validated MOFs,6 containing $10^{5}-10^{6}$ materials in total.
@@
In this work, we perform Quantum Chemistry calculations, on classical and on quantum computing hardware, in the form of the superconductor-based quantum computers available in the IBM Quantum platform, to estimate the binding potential of a set of representative molecules in the atmosphere (CO2, H2O and N2) within a simplified model system. We limited the search space for a carbon capture material by focusing on the X-MOF-74 family (where X=Mg+2, Ni+2, Fe+2, Co+2, Zn+2, Mn+2, Cu+2) as a template for exploring the chemical diversity in MOFs. The MOF-74 family is known for the presence of coordinatively unsaturated metal ions (open metal sites) that can form a bond with adsorbate molecules, greatly influencing the adsorption strength and selectivity, especially at low pressures. The formation of chemical bonds with the adsorbate molecule is a phenomenon known as chemisorption and determines, to a great extent, the desorption energy cost.kumar2015direct, decoste2013effect
@@ 1.1.1 Species selection
Given the relevance of the metal ions to DAC applications of MOFs, we chose to focus on Mg+2metal ion and its pairings with all the three provided gas molecules (H2O, CO2, and N2), amounting to six distinct systems. In short, Mg+2 + H2O should be the simplest system as it contains the least number of electrons.
@@ 1.1.2 Simulation methods
We chose PySCF,7, 8 as the driver for performing classical benchmarks. We performed two types of classical benchmarks with PySCF: Hartree-Fock (HF) and Density Functional Theory (DFT). The chosen basis set was 6-31g* and was kept consistent across all calculations, including for the quantum calculations. For DFT we chose the B3LYP functional. This particular choice of parameters was motivated by the methodology employed to compute 3D structures (SD-file) of the H2O, CO2, and N2 molecules found in the NIST Chemistry WebBook, from which we also took the molecular structures used in the calculations. 9
1.1.3 Considerations for Simulations on Quantum Computers
@@
  • A standard VQE algorithm with an application-specific, compute-heavy UCCSD ansatz,14 to be treated as the quantum benchmark, as it provides the best possible result for this kind of application. Unfortunately, it is challenging to use this approach on current quantum hardware, due to its complexity.
  • A standard VQE algorithm with a general-purpose, hardware-efficient heuristic ansatz tilly2022variational that can run on current quantum hardware. In our case, we determined that an ansatz with RY and RZ rotations, which we call RYRZ, converges to the intended result despite requiring more iterations.
  • A modified ADAPT-VQE algorithm15 that utilizes qubit operators in the operator pool. This method is called qubit-ADAPT-VQE and has the advantage of giving similar results to UCCSD while having much lower requirements on the quantum hardware.16
@@ 1.1.4 Active Space Selection
To limit the computational cost for the VQE calculations, we performed a suitable active space selection. We used a known method implemented within PySCF called atomic valence active space (AVAS). AVAS is an automated method to construct an active space for electronic structure calculations in strongly correlated chemical systems, where single-determinant wave functions are inadequate. It focuses on identifying specific valence atomic orbitals, such as transition metal d orbitals or those involved in bond dissociation, as a basis for defining the active space. The approach leverages density matrix embedding theory (DMET) to determine the minimal set of orbitals required to capture the essential electronic features, aiming for an active space no more than twice the size of the initially chosen orbitals. This method does not differentiate between “fragment” and “bath” orbitals, instead maintaining the occupied and virtual character of active orbitals, facilitating a natural truncation process to minimize active space size while capturing critical determinants.17
@@ 2.1 Convergence studies
The qubit-ADAPT-VQE ansatz accomplishes fast convergence while remaining hardware efficient, as we can see from Table 2 comparing the sizes of the circuits between methods with the largest circuit that qubit-ADAPT-VQE uses. These circuit sizes are expressed in terms of the gates that IBM quantum hardware can execute and were built using the transpilation procedure from Qiskit. The most important metric here is the number of CNOT gates as those have error rates that are orders of magnitude higher than the single-qubit gates.
Table 2: Comparing the circuit sizes generated by different ansatze. qubit-ADAPT-VQE manages to get satisfactory results while using the shallowest circuits. Circuits are described in terms of the number of gate operations and depth required to run on IBM quantum hardware.
@@ 2.2 Mg+2 systems
Due to the excellent cost-benefit obtained from the qubit-ADAPT-VQE ansatz in the simulations, we selected this as our best candidate to run the calculations on IBM Quantum hardware. However, to run these calculations on the available quantum hardware, we had to further reduce our active space to three electrons and four molecular orbitals (3,4), which only required 4 qubits. Despite nominally supporting higher qubit counts, current hardware are not suitable for reliably running circuits of the required depth for this application due to a lack of fault tolerance. Again, the molecular orbitals were chosen around the HOMO-LUMO level and ordered by energy. We set the limit on the number of shots to 1024, which takes from five to eight hours depending on the separation distance.
The 4-qubit active space limitation led us to focus only on the Mg+2 systems for the VQE calculations on quantum hardware which we refer to as hardware experiments. This active space corresponds to $\approx 5\%$ of the energy contributions to the ground state, therefore we saw no point in simulating systems with a higher electron count as it would mean an even lower representation of the results obtained. We used Qiskit Runtime and ibmq_montreal to run all our quantum hardware experiments, and we employed error mitigation techniques such as Twirled Readout Error Extinction (T-REx).19.
The PES was built by computing the energy for different distances between metal ions and molecules along the bond axis. For all simulations and hardware experiments, we evaluate the energy at distances between 0.3 and 3.6 Å, in increments of 0.3 Å.
@@
We simulated Mg+2 binding with all three gas molecules using the three VQE methods mentioned before: RYRZ-VQE, UCCSD-VQE, and qubit-ADAPT-VQE, on a (6,7) active space. We evaluated their accuracy by comparing their PES to that obtained via direct diagonalization of the Hamiltonian with NumPy.
@@
Based on the accuracy of qubit-ADAPT-VQE results in Figure 4, and the superior performance over UCCSD-VQE demonstrated in Figure 3, Table 1 and Table 2, we confirm the choice of qubit-ADAPT-VQE for running VQE on quantum hardware. We limit ourselves to the smallest possible system, namely, Mg+2 + H2O for the reasons explained previously. We ran qubit-ADAPT-VQEwith an active space of (3,4) to comply with the limitations of current quantum hardware.
The results, shown in Figure 5, also provide a comparison to a quantum hardware calculation. The Morse potential fit to the hardware result matches quite closely those of the direct diagonalization via NumPy and simulated VQE, despite the level of noise that is inherent to calculations on current quantum hardware. In this hardware experiment, we used 1024 shots and employed the T-REx noise mitigation technique which proved sufficient to handle the noise profile of the quantum device. Increasing the number of shots did not improve the accuracy of the experimental values, and using more robust error mitigation strategies, such as Zero Noise Extrapolation, would increase the already large variance in the results.
@@
We extracted observables such as the equilibrium distance and the binding energy from the adjusted Morse potentials to the Mg+2 + H2O PES curves. Although we cannot claim to have achieved agreement to within chemical accuracy ($\sim 1\ \text{mHa}$), in Table 3 we observe an excellent match between qubit-ADAPT-VQE and NumPy, and a reasonably good match with ibmq_montreal, considering the higher uncertainties caused by noise in the quantum hardware.
@@
In the following, we compare the potential energy surfaces computed using classical methods from PySCF, such as HF and DFT, to the best-in-class simulated VQE method, namely, qubit-ADAPT-VQE in a (6,7) active space. In Figure 6, we show the agreement between quantum methods and their classical benchmarks. In particular, for Mg+2 + H2O, we also include in Figure 6(a) the comparison to the quantum hardware experiments running qubit-ADAPT-VQE with ibmq_montreal in a (3,4) active space.
Overall, the simulated qubit-ADAPT-VQE results match the HF benchmark for all Mg+2-based systems studied, with the DFT benchmark coming quite close. The PES obtained by running in the quantum hardware, despite its inherent noise, is also in good agreement with the classical benchmarks. This conclusion is supported not only by the visual inspection of Figure 6 but also by a comparison of the numeric results of the observables extracted from the PES fit in Table 4. Here, the Binding Energy values all agree within the numeric uncertainty, and while the equilibrium distances between qubit-ADAPT-VQE and HF concur, they are found to be 2-3% higher than what is predicted by DFT.
It is worth noting the high uncertainties in our hardware results that are due to application of error mitigation techniques. While accuracy increases we typically observe a precision loss. Considering the high uncertainties observed in our results, we decided not to use advanced error mitigation techniques that would increase those uncertainties even further.
@@
Table 4: Comparison of Equilibrium Distances and Binding Energies for Mg+2-based systems obtained from the fitted Morse potential across different methods, both quantum (simulations and hardware experiments) and classical. The Binding Energy values all agree to within the numeric uncertainty. The equilibrium distances agree between qubit-ADAPT-VQE and HF, but are 2-3% higher than what is predicted by DFT.
@@ 2.3 Conclusions
The authors thank Max Rossmannek (IBM Research – Zurich) for providing help in implementing qubit-ADAPT-VQE and the broader Qiskit community for giving support. We thank Vidushi Sharma (IBM Research) for her support in the interpretation of the classical benchmark methods. We thank Felipe Lopes de Oliveira (IBM Research) for suggesting the Morse potential fit as a means to extract observables. We thank Ivan S. Oliveira (CBPF) and Alexandre Martins de Souza (CBPF) for insightful discussions on quantum computing and their support. The authors would like to acknowledge the support of Alexandre Pfeifer and Bruno Flach (IBM Research).